A barge takes 3 hours to go(at a constant rate) downstream 27 miles with a current of 3 miles per hour. If the barge's engines are set up at the same pace, find the time of the return trip against the current.
I am completely lost, and do not even know how to set up the problem.

Answer by scott8148(6628)

r = d / t

b + c = 27 / 3 ___ b + 3 = 9 ___ b = 6 ___ this is the barge's speed in still water

t = d / r

t = 27 / (6 - 3) ___ t = 9 ___ the current works against the barge going upstream